OpenStudy (anonymous):

the area of a rectangular wall of barn is 160 square feet. its length is 4 feet longer than twice its width. find the width of the wall of the barn

OpenStudy (anonymous):

(4ft+2w)w =160  (2w+4)(w)=160 2w^2+4w=160 \[2w^2+4w-160=0\] then factor to get : 2(w-8)(w+10)=0 set each equal to 0: w-8=0 w+10=0 w=8 w=-10 cant be negative so w=8ft

jhonyy9 (jhonyy9):

so than you need just substitute in formule the value of width what is 8 and check it than will get for result of area 160 ok ?

jhonyy9 (jhonyy9):

4 feet longer than twice its width - so than you know that the width is 8 than how many will be the length of ... ?

jhonyy9 (jhonyy9):

twice its width mean 2*8=16 so 4 feet longer than 16 = 16+4 =20 so than the length will be 20 and the width equal 8 ok ?

OpenStudy (anonymous):

yes whatever w is you substitute 8 in for that for the length it would be 2(8)+4= 20 20*8=160. so ya you are right

jhonyy9 (jhonyy9):

yes hope you understand it sure right good luck bye